Hello..... I need help to determine if the graphs of the given equation are parallel. I am going to be honest i dont have a clue on what i am doing
y=4x+3
3y+4x=8

Found 2 solutions by rfer, TimothyLamb:
Answer by rfer(16322)

y=4x+3
------------
3y+4x=8
3y=-4x+8
y=-4/3x+8/3
----
put both in y intercept form, the slopes are different so
they are not parallel.

Answer by TimothyLamb(4379)

put the equation is slope-intercept form: y = mx + b
y = 4x + 3
3y + 4x = 8
---
y = 4x + 3
this equation is already in slope-intercept form: y = mx + b
m = 4
b = 3
---
3y + 4x = 8
this equation is not in slope-intercept form: y = mx + b
(3/3)y + (4/3)x = 8/3
y = (-4/3)x + 8/3
but now it is:
m = -4/3
b = 8/3
---
two lines are parallel only if they have the same slope
---
Answer:
these lines do not have the same slope so they are not parallel
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php